# Vacuum amplitudes and time-like causal unitary in the loop-tree duality

### v1

@@ V Numerical implementation
If $p_{1}$ and $p_{2}$ are the momenta of the quark and the antiquark, respectively, in the final-state with two particles, and $p_{1}^{\prime}$, $p_{2}^{\prime}$ and $p_{3}^{\prime}$ are the momenta of the quark, antiquark and gluon in the final state with three particles, the following mappings apply
@@
The phase-space energy function $\widetilde{\Delta}_{35\bar{6}78}$ in Eq. (35) and its symmetric counterpart $\widetilde{\Delta}_{24\bar{6}78}$ introduce a dependence in the $\bm{\ell}_{4}$ loop momentum, and therefore two additional independent integration variables. We should also consider final-states with four external particles and the mappings
@@
$q_{7,0}^{(+)}=\sqrt{\bm{\ell}_{4}^{2}+m^{2}-\imath 0}~{},$
$q_{8,0}^{(+)}=\sqrt{(\bm{\ell}_{4}+\bm{\ell}_{12})^{2}+m^{2}-\imath 0}~{}.$ (49)
The numerical implementation is quite stable, in particular due to the absence of unitary threshold singularities. The numerical results obtained for the NLO contribution to the decay rates, normalised to the LO total decay rate, are presented in Fig. 3, where they are compared with state-of-the-art DREG analytical expressions. The agreement shown is excellent. Only $H\to q\bar{q}$ exhibits a dependence on the renormalisation scale $\mu_{\rm UV}$ because the vector current of the photon decay is UV protected and the scalar decay is UV finite at NLO.
In Fig. 4 and Fig. 5, we illustrate the integrand behaviour of the different phase-space residues and their sum across a unitary threshold singularity and in collinear configurations, respectively. We can clearly observe how the different phase-space residues are singular, but their sum provides a flat integrand. The phase-space residues shown in Fig. 4 and Fig. 5 correspond to $H\to q\bar{q}$. Similar results are obtained for the other two processes.
The local cancellation of collinear singularities at NNLO is illustrated in Fig. 6 for the toy model described in Section IV. We have fixed the angular variable that defines the collinearity of particles $7$ and $8$ to $v_{4}=10^{-2}$. We can observe that for large values of the other angular variable, $v$, the cancellation occurs mostly between ${\cal A}_{\rm D}^{(4,\Phi)}(1356)$ and ${\cal A}_{\rm D}^{(4,\Phi)}(35678)$, which corresponds to the local cancellation of the double-collinear singularity at $\lambda_{\bar{1}78}\to 0$ according to Eq. (40). At very small angles, $v<v_{4}$, instead ${\cal A}_{\rm D}^{(4,\Phi)}(456)$ becomes much more singular than ${\cal A}_{\rm D}^{(4,\Phi)}(35678)$, reflecting the dominance of the emergent triple-collinear configuration, and the participation of the three phase-space residues is necessary for the local cancellation of the overlapping collinear singularities, according to Eq. (41).
@@
Figure 4: Unintegrated decay rate across a unitary threshold singularity as a function of the angular variable. The sum over the one-loop phase-space residues is flat.
VI Conclusions
We have presented the first proof-of-concept implementation of LTD causal unitary [1] to decay processes at higher perturbative orders. The processes chosen are such that all the benefits of the method are clearly visible in a lightweight manner. Specifically, the central and novel ingredient of LTD causal unitary is the use of a multiloop vacuum amplitude in the LTD representation as a kernel of all the final states contributing to the decay process. The LTD representation of the vacuum amplitude, which is manifestly causal, coherently encodes these final states as residues on causal propagators that depend on linear combinations of the on-shell energies of the internal propagators, once certain on-shell energies identified with the incoming particles are analytically continued to negative values.
@@ VI Conclusions
Concerning unitary threshold singularities, which typically represent an overwhelming obstacle in numerical implementations, LTD causal unitary provides an efficient solution leading to flat integrands across thresholds, as the potential singularities match between phase-space residues with the same number of final-state particles. This property, which is a natural consequence of using a vacuum amplitude as a kernel, represents a clear advantage over other methods.
The local cancellation of singularities at NLO and NNLO has been illustrated with selected decay processes. The total decay rates at NLO have been obtained with LTD causal unitary and have been compared with state-of-the-art DREG analytical expressions, showing a perfect agreement. The transition from a massive to a massless implementation, as already observed in [20], is smooth. The numerical implementation in LTD causal unitary is quite stable and leads to accurate results with minimal CPU resources.
The results presented in this paper constitute a solid confirmation of the unique capabilities and advantages of LTD causal unitary at higher perturbative orders, in addition to successfully addressing the original motivation and path initiated with the seminal works on the loop-tree duality [74, 75, 76]. A proof-of-concept implementation for scattering processes that require a local subtraction of initial-state collinear singularities, as well as more complex and well-motivated phenomenological analysis with LTD causal unitary for scattering and decay processes at high-energy colliders, will be presented in forthcoming publications.
@@ Acknowledgements
This work is dedicated to the memory of Stefano Catani, who always inspired us in the development of this research line, and very touching during the Workshop Theory Challenges in the Precision Era of the Large Hadron Collider held at the Galileo Galilei Institute (Firenze) in September 2023. This work is supported by the Spanish Government (Agencia Estatal de Investigación MCIN /AEI/10.13039/501100011033) Grants No. PID2020-114473GB-I00 and PID2022-141910NB-I00, and Generalitat Valenciana Grant No. PROMETEO/2021/071. AERO is supported by the Spanish Government (PRE2018-085925). DFRE and JML are supported by Generalitat Valenciana (CIGRIS/2022/145 and ACIF/2021/219). PKD is supported by European Commission MSCA Action COLLINEAR-FRACTURE, Grant Agreement No. 101108573. SRU and RJHP are funded by CONAHCyT through Project No. 320856 (Paradigmas y Controversias de la Ciencia 2022), Ciencia de Frontera 2021-2042 and Sistema Nacional de Investigadores. LC is supported by Generalitat Valenciana GenT Excellence Programme (CIDEGENT/2020/011) and ILINK22045. GS is partially supported by EU Horizon 2020 research and innovation programme STRONG-2020 project under Grant Agreement No. 824093 and H2020-MSCA-COFUND USAL4EXCELLENCE-PROOPI-391 project under Grant Agreement No 101034371. WJT is supported by the Leverhulme Trust, LIP-2021-01.
@@ Appendix A The phase space
$d\Phi_{m}$ $=$ $\mu^{d-4}\,(2\pi)^{d-1}\,\delta^{(d-1)}\left(\sum_{i=1}^{m}\mathbf{p}_{i}-\mathbf{p}_{ab}\right)\,$ (53)
$\times$ $\frac{\widetilde{\Delta}_{1\cdots m\bar{a}\bar{b}}}{x_{1\cdots m}}\,d\Phi_{{\bf p}_{1},\ldots,{\bf p}_{n}}~{},$

### v2

@@ 5 Numerical implementation
If $p_{1}$ and $p_{2}$ are the momenta of the quark and the antiquark, respectively, in the final-state with two particles, and $p_{1}^{\prime}$, $p_{2}^{\prime}$ and $p_{3}^{\prime}$ are the momenta of the quark, antiquark and gluon in the final state with three particles, the following linear mappings apply
@@
Figure 4: Unintegrated decay rate across a threshold singularity as a function of the angular variable. The sum over the one-loop phase-space residues is flat.
The phase-space energy function $\widetilde{\Delta}_{35\bar{6}78}$ in Eq. (35) and its symmetric counterpart $\widetilde{\Delta}_{24\bar{6}78}$ introduce a dependence in the $\bm{\ell}_{4}$ loop momentum, and therefore two additional independent integration variables. We should also consider final-states with four external particles and the linear mappings
@@
$q_{7,0}^{(+)}=\sqrt{\bm{\ell}_{4}^{2}+m^{2}-\imath 0}~{},\qquad q_{8,0}^{(+)}=\sqrt{(\bm{\ell}_{4}+\bm{\ell}_{12})^{2}+m^{2}-\imath 0}~{}.$ (49)
The numerical implementation is quite stable, in particular due to the absence of threshold singularities. The numerical results obtained for the NLO contribution to the decay rates, as a function of the final-state mass and normalised to the LO total decay rate, are shown in Fig. 3, where they are compared with the state-of-the-art DREG analytical expressions (see Appendix B). The corresponding numerical values are presented in Table 1. We have used a private version of VEGAS Lepage:1977sw which achieves a precision of $10^{-3}-10^{-4}$ with $10$ iterations of $10^{7}$ calls. This is a few seconds per point on a personal laptop. Only $H\to q\bar{q}$ exhibits a dependence on the renormalisation scale $\mu_{\rm UV}$ because the vector current of the photon decay is UV protected and the scalar decay is UV finite at NLO. Similar results have been obtained by using Quantum Fourier Iterative Amplitude Estimation (QFIAE) deLejarza:2023qxk ; deLejarza:2024pgk . This is a quantum algorithm for numerical integration of multidimensional functions that trains a quantum neural network to decompose the target function into its Fourier components, and then takes advantage of the fact that trigonometric functions are efficiently integrable in a quantum approach. We refer the interested reader to Ref. deLejarza:2024scm .
In Fig. 4 and Fig. 5, we illustrate the integrand behaviour of the different phase-space residues and their sum across a threshold singularity and in collinear configurations, respectively. We can clearly observe how the different phase-space residues are singular, but their sum provides a flat integrand. The phase-space residues shown in Fig. 4 and Fig. 5 correspond to $H\to q\bar{q}$. Similar results are obtained for the other two processes.
The local cancellation of collinear singularities at NNLO is illustrated in Fig. 6 for the toy model described in Section 4. We have fixed the angular variable that defines the collinearity of particles $7$ and $8$ to $v_{4}=10^{-2}$. We can observe that for large values of the other angular variable, $v$, the cancellation occurs mostly between ${\cal A}_{\rm D}^{(4,\Phi)}(1356)$ and ${\cal A}_{\rm D}^{(4,\Phi)}(35678)$, which corresponds to the local cancellation of the double-collinear singularity at $\lambda_{\bar{1}78}\to 0$ according to Eq. (40). At very small angles, $v<v_{4}$, instead ${\cal A}_{\rm D}^{(4,\Phi)}(456)$ becomes much more singular than ${\cal A}_{\rm D}^{(4,\Phi)}(35678)$, reflecting the dominance of the emergent triple-collinear configuration, and the participation of the three phase-space residues is necessary for the local cancellation of the overlapping collinear singularities, according to Eq. (41).
@@
6 Conclusions
We have presented the first proof-of-concept implementation of LTD causal unitary Ramirez-Uribe:2024rjg to decay processes at higher perturbative orders. The processes chosen are such that all the benefits of the method are clearly visible in a lightweight manner. Specifically, the central and novel ingredient of LTD causal unitary is the use of a multiloop vacuum amplitude in the LTD representation as a kernel of all the states contributing to the decay process. The LTD representation of the vacuum amplitude, which is manifestly causal, coherently encodes the final states as residues on causal propagators that depend on linear combinations of the on-shell energies of the internal propagators, once certain on-shell energies identified with the incoming particles are analytically continued to negative values.
@@ 6 Conclusions
Concerning threshold singularities, which typically represent an overwhelming obstacle in numerical implementations, LTD causal unitary provides an efficient solution leading to flat integrands across thresholds, as the potential singularities match between phase-space residues with the same number of final-state particles. This property, which is a natural consequence of using a vacuum amplitude as a kernel, represents a clear advantage.
The local cancellation of singularities at NLO and NNLO has been illustrated with selected decay processes. The total decay rates at NLO have been obtained with LTD causal unitary and have been compared with state-of-the-art DREG analytical expressions, showing a perfect agreement. The transition from a massive to a massless implementation, as first observed in Sborlini:2016hat , is smooth. The numerical implementation in LTD causal unitary is quite stable and leads to accurate results with minimal CPU resources.
The results presented in this paper constitute a solid confirmation of the unique capabilities and advantages of LTD causal unitary at higher perturbative orders, in addition to successfully addressing the original motivation and path initiated with the seminal works on the loop-tree duality Catani:2008xa ; Bierenbaum:2010cy ; Bierenbaum:2012th . A proof-of-concept implementation for scattering processes that require a local subtraction of initial-state collinear singularities, as well as more complex and well-motivated phenomenological analysis with LTD causal unitary for scattering and decay processes at high-energy colliders, will be presented in forthcoming publications.
@@ Acknowledgements
This work is dedicated to Stefano Catani, who always inspired us in the development of this research line. Part of this work was completed during the Workshop Theory Challenges in the Precision Era of the Large Hadron Collider held at the Galileo Galilei Institute (Firenze) in September 2023. This work is supported by the Spanish Government (Agencia Estatal de Investigación MCIN/ AEI/10.13039/501100011033) Grants No. PID2020-114473GB-I00, No. PID2022-141910NB-I00, No. PID2023-146220NB-I00 and No. CEX2023-001292-S, and Generalitat Valenciana Grant No. PROMETEO/2021/071. AERO is supported by the Spanish Government (PRE2018-085925). DFRE and JML are supported by Generalitat Valenciana (CIGRIS/2022/145 and ACIF/2021/219). PKD is supported by European Commission MSCA Action COLLINEAR-FRACTURE, Grant Agreement No. 101108573. SRU and RJHP are funded by CONAHCyT through Project No. 320856 (Paradigmas y Controversias de la Ciencia 2022), Ciencia de Frontera 2021-2042 and Sistema Nacional de Investigadores. LC is supported by Generalitat Valenciana GenT Excellence Programme (CIDEGENT/2020/011) and ILINK22045. GS is partially supported by EU Horizon 2020 research and innovation programme STRONG-2020 project under Grant Agreement No. 824093 and H2020-MSCA-COFUND USAL4EXCELLENCE-PROOPI-391 project under Grant Agreement No 101034371. WJT is supported by the Leverhulme Trust, LIP-2021-01.
@@ Appendix A The phase space
$d\Phi_{m}$ $=$ $\mu^{d-4}\,(2\pi)^{d-1}\,\delta^{(d-1)}\left(\sum_{i=1}^{m}\mathbf{p}_{i}-\mathbf{p}_{ab}\right)\,\frac{\widetilde{\Delta}_{1\cdots m\bar{a}\bar{b}}}{x_{1\cdots m}}\,d\Phi_{{\bf p}_{1},\ldots,{\bf p}_{n}}~{},$ (53)
@@
Appendix B DREG analytical expressions for NLO decay rates
The following expressions for the NLO predictions of the decay rates from DREG are adapted from Ref. Sborlini:2016hat . First, we define the function
$h_{p}=\frac{4}{\beta}\left[\mathrm{log}\left(x_{S}\right)\left(\mathrm{log}\left(1-x_{S}\right)+\mathrm{log}\left(1+x_{S}^{2}\right)\right)+2\mathrm{Li_{2}}\left(x_{S}\right)+\mathrm{Li_{2}}\left(x_{S}^{2}\right)\right]~{},$ (56)
where $x_{S}=(1-\beta)/(1+\beta)$, with $\beta=\sqrt{1-4m^{2}/s}$. We also define the functions
${\cal F}_{H}$ $=$ $(1+\beta^{2})h_{p}-2\mathrm{log}\left(\frac{m^{2}}{s}\right)-4\left(\frac{1-2\beta}{\beta}\mathrm{log}\left(x_{S}\right)+4\mathrm{log}\left(1-x_{S}\right)-5\right)~{},$
${\cal F}_{\gamma^{*}}$ $=$ $(1+\beta^{2})h_{p}-2\mathrm{log}\left(\frac{m^{2}}{s}\right)$ (57)
-- $4\left(\frac{2-7\beta}{2\beta}\mathrm{log}\left(x_{S}\right)+4\mathrm{log}\left(1-x_{S}\right)+3\mathrm{log}\left(1+x_{S}\right)-3\right)~{},$
and
${\cal G}_{H}=-3\left(1+\beta^{2}\right)+\frac{13\beta^{4}-18\beta^{2}-3}{2\beta}\mathrm{log}\left(x_{S}\right)~{},$
${\cal G}_{\gamma^{*}}=3\left(\beta^{2}-7\right)+\frac{7\beta^{4}-30\beta^{2}-9}{2\beta}\mathrm{log}\left(x_{S}\right)~{}.$ (58)
The coupling striped theoretical prediction for the decay rates at NLO, normalized to the LO, is
${\bf\Gamma}_{f}^{(1)}=\frac{g_{f}^{(0)}}{g_{f}^{(1)}}\frac{\Gamma_{f}^{(1)}}{\Gamma_{f}^{(0)}}~{},$ (59)
with
${\bf\Gamma}_{\Phi}^{(1)}=\frac{h_{p}}{16\pi^{2}}~{},$
${\bf\Gamma}_{H}^{(1)}=\frac{1}{16\pi^{2}}\left[6\mathrm{log}\left(\frac{s}{\mu_{\rm UV}^{2}}\right)+{\cal F}_{H}+\frac{{\cal G}_{H}}{2\beta^{2}}\right]~{},$
${\bf\Gamma}_{\gamma^{*}}^{(1)}=\frac{1}{16\pi^{2}}\left[{\cal F}_{\gamma^{*}}+\frac{{\cal G}_{\gamma^{*}}}{3-\beta^{2}}\right]~{}.$ (60)
Only the theoretical prediction for the Higgs boson decay has a dependence on the renormalisation scale $\mu_{\rm UV}$.